# QOS: A Quantum Operating System

### v1

@@
Abstract.
We introduce the Quantum Operating System (QOS), a unified system stack for managing quantum resources while mitigating their inherent limitations, namely their limited and noisy qubits, (temporal and spatial) heterogeneities, and load imbalance. QOS features the QOS compiler—a modular and composable compiler for analyzing and optimizing quantum applications to run on small and noisy quantum devices with high performance and configurable overheads. For scalable execution of the optimized applications, we propose the QOS runtime—an efficient quantum resource management system that multi-programs and schedules the applications across space and time while achieving high system utilization, low waiting times, and high-quality results.
@@ Abstract.
We evaluate QOS on real quantum devices hosted by IBM, using 7000 real quantum runs of more than 70.000 benchmark instances. We show that the QOS compiler achieves 2.6–456.5$\times$ higher quality results, while the QOS runtime further improves the quality by 1.15–9.6$\times$ and reduces the waiting times by up to 5$\times$ while sacrificing only 1–3% of results quality (or fidelity).
1. Introduction
@@
Quantum Cloud Computing. Quantum computing promises to solve computationally intractable problems with classical computers (Arute et al., 2019; Farhi et al., 2014). Thanks to remarkable technological advances in materials science and engineering (Siddiqi, 2021; Hanneke et al., 2010), quantum hardware has become a reality in the form of quantum processing units (QPUs) that consist of quantum bits (qubits) (ibmQuantum, [n. d.]). Interestingly, QPUs are now readily available in a quantum-as-a-service fashion offered by all major cloud providers (ibmQuantum, [n. d.]; awsQuantum, [n. d.]; googleQuantum, [n. d.]; azurequantum, [n. d.]).
While quantum hardware is now a reality, the associated quantum software systems are rudimentary. These QPUs face classical OS challenges that our systems community has tackled in the past, including scalability, performance, efficiency, faults (a.k.a. errors), scheduling, and utilization (Bravyi et al., 2022). Unfortunately, no operating system exists to tackle these challenges holistically for modern quantum hardware.
Fundamental Challenges of QPUs. A natural tendency would be to treat these QPUs as yet another accelerator class (e.g., GPU, TPU, FPGA) and manage them as accelerator as-a-service to offload compute-intensive tasks. We argue that this approach might be sub-optimal or even flawed!
@@
The reason is that QPUs present fundamentally unique hardware-level challenges that the systems community has not considered and cannot be directly mapped to classical accelerator-oriented computing (we empirically detail these hardware challenges in §$\S$ 3). In particular, QPUs operate in the NISQ-fashion (Noisy Intermediate-Scale Quantum (Preskill, 2018)), leading to a non-deterministic computing platform, where even two QPUs with identical qubits exhibit completely different behaviors across space and time (Ravi et al., 2021a; Patel et al., 2020).
@@
More specifically, QPUs are inherently noisy and small in computational capacity (Preskill, 2018), which limits the size of the problems they can solve. Second, the degree of noise differs across QPUs, even of identical architecture and model, making it difficult to decide which QPUs should execute a quantum program without compromising performance (Ravi et al., 2022b). In addition, we can not trivially multi-program multiple quantum programs on the same QPU to increase utilization since QPU qubits can interfere with each other in undesirable and unpredictable ways (Murali et al., 2020), severely degrading performance (Liu and Dou, 2021). Finally, it is generally impossible to save or copy a quantum program during execution (Park, 1970), which further limits scheduling opportunities for preemption or resource sharing in general.
State-of-the-Art of Quantum Software Systems. The current state of software can be roughly compared to IBM mainframe batch OSes from the 60s, where the QPUs are managed through rudimentary interfaces. Researchers have proposed specialized approaches to address some of the aforementioned OS and QPU challenges individually, for instance, performance (Tang et al., 2021a), multi-programming (Das et al., 2019), or scheduling (Ravi et al., 2021b). Unfortunately, these proposed approaches are designed to solve an individual issue, which prevents them from being composed together or with other OS mechanisms to create a holistic software stack. To leverage quantum computing practically, we must address the key challenge of combining such mechanisms in a unified software stack for quantum computing.
Novelty. However, designing a unified system stack that supports general OS abstractions while addressing the QPU challenges is not trivial. The system should support cross-stack software mechanisms, from the compiler level for quantum program optimization to the runtime level for QPU resource management. More specifically, we require a modular and extensible compiler infrastructure for increasing execution quality, multi-programming for increased utilization, and scheduling for load balancing, all in the presence of QPU noise and heterogeneities. This way, the system can achieve the cloud users’ goals, i.e., high-quality quantum computation and low waiting times, and the quantum cloud operator’s goals, i.e., QPU resource efficiency and scalability.
@@
Figure 1. Example of a typical quantum algorithm (§ 2.1) (a) Input graph for max-cut. (b) The quantum circuit encoding the formulation of max-cut for the graph. (c) The result of circuit execution is a probability distribution of bitstrings. (d) The result of (c) is interpreted as a max-cut between vertices $\{0,1,4\}$ and $\{2,3,5\}$.
QOS: A Unified System Stack for Quantum Computing. We propose QOS, an end-to-end system for holistically tackling quantum computing challenges. QOS provides a unified architecture for supporting compiler and OS mechanisms with pluggable and configurable policies. In QOS, we implement such policies to achieve the aforementioned users’ and operator’s goals. To achieve this, QOS builds on a unified abstraction and comprises two main components:
  • The Qernel Abstraction: We introduce the Qernel abstraction that acts as a common denominator for the QOS mechanisms to apply their policies (§ 5.1). A Qernel contains the Qernel intermediate representation (QIR) and static and dynamic properties, leveraged by the QOS components to apply their policies.
  • QOS Compiler: We introduce the QOS compiler (§ 5, 6, 7), an extensible and modular compiler workflow that leverages the QIR and static properties to optimize quantum programs for increased execution quality.
  • QOS Runtime: We present the QOS runtime (§ 8), a scalable system for QPU resource efficiency. The system offers automated QPU selection to abstract heterogeneity away, multi-programming to increase QPU utilization, and load-aware scheduling to achieve low waiting times while maintaining high execution quality.
Figure 2. Technical Foundations (§ 2.2) (a) The quantum circuit of Figure 1. (b) The physical layout of an IBM Falcon QPU. (c) The transpiled circuit with the QPU’s noise sources.
@@
We implement QOS in Python by building on the Qiskit framework (Qiskit contributors, 2023). We evaluate QOS on IBM’s 27-qubit QPUs (ibmQuantum, [n. d.]), using a dataset of more than 7000 quantum runs and 70.000 state-of-the-art quantum benchmark instances used in popular quantum algorithms (Tomesh et al., 2022; Quetschlich et al., 2022; Li et al., 2021). Our evaluation shows that the QOS compiler improves the quantum program properties by 51% on average, which leads to 2.6–456.5$\times$ improvement in the quality of the results, depending on the problem size (§ 9.2). The QOS runtime increases the quality of the results by 1.15–9.6$\times$ for the same target utilization (§ 9.4) and reduces the waiting times by 5$\times$ while sacrificing at most 3% of the quality of the results (§ 9.5).
2. Background
2.1. Quantum Computing 101: An Example
Let us understand the basics of quantum computing using the classic max-cut problem. This simple combinatorial optimization problem is expressed in the quantum world as the Quantum Approximate Optimization Algorithm (QAOA) (Farhi et al., 2014). Figure 1 shows a high-level example of how QAOA solves a max-cut problem of the input graph of (a). To solve it, the problem must be first encoded as a quantum circuit (Figure 1 (b)), which consists of quantum bits (qubits) and quantum gates that exhibit quantum mechanical properties. Here, we use as many qubits as the number of nodes of the input graph, where each qubit $q_{i}$ corresponds to a graph vertex $i$. To change the state of the qubits, we apply quantum gates over time, from left to right. There are two types of gates: 1-qubit gates (e.g., NOT gate) and 2-qubit gates (e.g., XOR gate). Finally, at the end of the circuit, we measure each qubit to read its value (0 or 1), which gives bitstrings as output.
Unlike classical circuits, which operate deterministically, quantum circuits are inherently probabilistic. The reason is that qubits exhibit quantum mechanical properties, such as superposition. In the superposition state, the qubit is not 0 or 1, but it is both simultaneously (recall Schrodinger’s cat experiment (schrondinger, [n. d.])). Therefore, quantum gates also have probabilistic effects; we can’t know the result until the final measurements (i.e., open the box and check the cat’s state). To obtain a meaningful result, we execute the circuit in many trials (“shots”), with each trial providing a specific bitstring from the qubit measurement. The solution of the quantum calculation is, therefore, a probability distribution over all possible bitstrings of the measured qubits (Figure 1 (c)).
@@ 2.1. Quantum Computing 101: An Example
Figure 3. (a) Challenge #1, Fidelity (§ 3.1). Impact of the number of qubits (circuit size) on fidelity. There is an average 98.9% reduction in fidelity from 4 to 24 qubits. (b) Challenge #2, Spatial heterogeneity (§ 3.2) Fidelity of a 12-qubit GHZ circuit on different IBM QPUs. There is a 38% fidelity difference from best to worst QPU.
In our example, the result of the final execution of the quantum circuit gives a probability distribution that represents the solutions of the max-cut problem. High probability maps to the solution, while low ($\sim 0$) does not represent a solution. Figure 1 (d) shows a solution for our example. It corresponds to the bitstring with the highest probability, 110010, which means that we have measured $1$ for the qubits $q_{0}$, $q_{1}$, and $q_{4}$; therefore, a partition contains vertices $\{0,1,4\}$.
2.2. Technical Foundations
@@
Execution Model. The technology and engineering required to build QPUs renders them an expensive resource, thus, QPUs are mainly offered in the cloud as a quantum-as-a-service model (ibmQuantum, [n. d.]; googleQuantum, [n. d.]; awsQuantum, [n. d.]). To run quantum programs, users typically write circuit-level code (Figure 2 (a)), which then transpile on the QPU to make it executable, send it to the cloud for execution, and finally get the results back. Specifically, the transpilation process performs three key steps: (1) converting the gates of the circuit to the native gate set of the QPU, (2) mapping the logical qubits of the circuit to the physical qubits of the QPU, (3) routing the qubits to the physical qubits with restrictive connectivity by inserting SWAP gates. Figure 2 (b) shows the physical layout of an IBM Falcon QPU. Vertices are the physical qubits, and the edges capture their connectivity, i.e., between which qubits we can apply 2-qubit gates. Figure 2 (c) shows the physical circuit after transpilation with the QPU’s noise characteristics, which we detail next.
@@
QPU Characteristics. Today’s QPUs are described as noisy intermediate-scale quantum (NISQ) devices (Preskill, 2018) since they exhibit low qubit numbers (e.g., up to a few 100s (ibmQuantum, [n. d.])) and are susceptible to hardware and environmental noise. Specifically, when measuring a qubit, there is a chance to read the opposite value, and when applying gates, there is a chance the gate performs a wrong operation (google-nisq-properties, [n. d.]). On top of that, when qubits are left idle (no gates applied) for more than a few hundred microseconds, the superposition decoheres to the |$\ket{0}$ state (Klimov et al., 2018), similar to resetting a register to 00. Lastly, qubits destructively interfere with each other via crosstalk effects (Cross et al., 2019). Figure 2 (c) shows qubits $Q_{2}$ and $Q_{3}$ that influence each other via crosstalk, noisy gates, qubit $Q_{5}$ that is left idle for long enough to decohere, and noisy measurements.
@@
QPU Heterogeneity. Additionally, QPUs are vastly heterogeneous across space and time, unlike classical accelerators. Across space, QPUs vary in terms of technology, e.g., superconducting qubits (googleQuantum, [n. d.]; ibmQuantum, [n. d.]) or trapped ions (Häffner et al., 2008), architectures of the same technology, e.g., Falcon or Osprey superconducting QPUs (ibmQuantum, [n. d.]), and noise properties even for the same architecture (google-nisq-properties, [n. d.]), e.g., two identical QPUs exhibit different noise errors, etc. Across time, the QPUs are calibrated regularly to maintain their performance (Wittler et al., 2021; Tornow et al., 2022; ibmq-calibration, [n. d.]), a process that generates calibration data. These data quantify the noise errors, and change after each calibration cycle unpredictably.
@@
Execution Quality. Lastly, to measure the quality of a circuit execution on NISQ QPUs, we use the fidelity metric (fidelity-qiskit, [n. d.]), which measures the similarity between the noisy probability distribution and the ideal probability distribution that noiseless, ideal QPUs can obtain. Fidelity is a number in the $[0,1]$ range, where a higher fidelity means a better quality result.
3. Motivation and Key Ideas

### v2

@@
Abstract
Quantum computers face challenges due to hardware constraints, noise errors, and heterogeneity, and face fundamental design tradeoffs between key performance metrics such as quantum fidelity and system utilization. This substantially complicates managing quantum resources to scale the size and number of quantum algorithms that can be executed reliably in a given time.
We introduce QOS, a cloud operating system for managing quantum resources while mitigating their inherent limitations and balancing the design tradeoffs of quantum computing. QOS exposes a hardware-agnostic API for transparent quantum job execution, mitigates hardware errors, and systematically multi-programs and schedules the jobs across space and time to achieve high quantum fidelity in a resource-efficient manner. To achieve this, it leverages two key insights: First, to maximize utilization and minimize fidelity loss, some jobs are more compatible than others for multi-programming on the same quantum computer. Second, sacrificing minimal fidelity can significantly reduce job waiting times.
@@ Abstract
We evaluate QOS on real quantum devices hosted by IBM, using 7000 real quantum runs of more than 70.000 benchmark instances. We show that the QOS achieves 2.6–456.5$\times$ higher fidelity, increases resource utilization by up to 9.6$\times$, and reduces waiting times by up to 5$\times$ while sacrificing only 1–3% fidelity, on average, compared to the baselines.
1 Introduction
@@
Quantum computing promises to solve computationally intractable problems with classical computers [2, 12, 25, 68]. Thanks to remarkable technological advances in materials science and engineering, quantum hardware has become a reality in the form of quantum processing units (QPUs) [69, 27, 19]. Interestingly, QPUs are now readily available in a quantum-as-a-service fashion offered by all major cloud providers [32, 3, 24, 5].
@@
However, QPUs present fundamentally unique hardware-level challenges that cannot be directly mapped to classical accelerator-oriented computing (we empirically detail these hardware challenges in § 3). In particular, QPUs operate in the NISQ-fashion (Noisy Intermediate-Scale Quantum [57]), leading to strict hardware constraints and a non-deterministic computing platform [61, 50].
@@
More specifically, QPUs are inherently noisy and small in computational capacity, which limits the size of the problems they can solve [57]. Second, the degree of noise differs across QPUs and time, even of identical architecture and model, making it difficult to decide which QPUs should execute a quantum program with good performance [61]. In addition, we can not trivially multi-program multiple quantum jobs on the same QPU to increase utilization without them interfering with each other in undesirable and unpredictable ways [45], severely degrading performance [40]. Last, scheduling quantum jobs on heterogeneous noisy QPUs exhibits a fundamental tension between quantum performance metrics, such as fidelity [21], and classical performance metrics, such as utilization and load balance (i.e., resource efficiency).
Unfortunately, these problems are amplified by the current state of software, where QPUs are managed through rudimentary interfaces, despite the ever-increasing demand for these scarce resources [61, 50, 1]. Researchers have proposed specialized approaches to address some of the aforementioned OS and QPU challenges individually, for instance, fidelity without runtime support [4], basic multi-programming via FIFO or random job selection [16], or heuristics-based scheduling [62]. Unfortunately, these systems are designed in isolation, are tightly coupled to specific policies, and lack the architectural flexibility and interoperability needed for cross-stack coordination.
At the same time, in industry platforms like IBM Cloud  [32] and AWS Braket [3], multi-programming is not even supported, and users must improve fidelity and select QPUs manually. However, the noisy, scarce, and heterogeneous nature of QPUs naturally motivates users to select the best-performing QPUs, further amplifying the already high QPU load imbalance and queue waiting times [61, 50].
In total, there is still no cloud operating system that addresses quantum job and resource management in a unified and systematic way, i.e., an architecture that supports composable, cross-layer mechanisms designed specifically for the constraints of quantum computing. Such a system must integrate error mitigation techniques to improve fidelity, fidelity-aware performance estimation to handle the spatiotemporal variability of QPUs, multi-programming to increase utilization through compatibility-aware co-location of quantum jobs, and a scheduler that balances QPU load and reduces queue times—all under a single, unified stack.
To fill this gap, we propose QOS, a cloud operating system for holistically tackling quantum computing challenges through a modular, cross-layer architecture. At the core of QOS lies the Qernel abstraction—the common denominator that enables the seamless composition of diverse system mechanisms and facilitates cross-layer optimizations.
QOS abstracts away the underlying complexity of quantum resource management and systematically explores the associated tradeoffs of quantum computing. To achieve this, QOS exposes hardware-agnostic APIs and comprises four main components: (1) the error mitigator, a component that composes complementary techniques to increase fidelity by mitigating hardware noise, (2) the estimator, which predicts fidelity performance across heterogeneous QPUs to guide scheduling decisions, (3) the multi-programmer that bundles compatible jobs on the same QPU to improve QPU utilization while minimizing fidelity loss, and (4), the scheduler, which performs fidelity-aware, multi-objective job scheduling to reduce queueing latency and balance QPU load.
@@
Figure 1: Foundational example (§ 2.1) (a) Input graph to max-cut. (b) A quantum circuit encoding the max-cut formulation for the graph. (c) The execution result is a probability distribution of bitstrings. (d) The result is interpreted as a max-cut between vertices $\{a,d\}$ and $\{b,c,e\}$.
@@
We implement QOS in Python by building on the Qiskit framework [33]. We evaluate QOS on IBM’s 27-qubit QPUs [32], using a dataset of more than 7000 quantum runs and 70.000 state-of-the-art quantum benchmark instances used in popular quantum algorithms [76, 60, 38]. Our evaluation shows that the error mitigator improves execution fidelity by 2.6–456.5$\times$ on average, depending on the problem size (§ 9.2), the estimator correctly identifies high-fidelity QPUs (§ 9.3), the multi-programmer fidelity by 1.15–9.6$\times$ for a target utilization (§ 9.4), and the scheduler reduces the waiting times by 5$\times$ while sacrificing at most 3% of fidelity (§ 9.5).
Contributions. Our main contributions include:
  1. 1.
    QOS is the first quantum operating system to holistically address the challenges of quantum computing. Its modular architecture and Qernel abstraction enforce a clean separation of mechanism and policy, enabling flexible integration of evolving techniques without modifying system interfaces.
  2. 2.
    QOS enables both cross-layer and intra-layer optimizations through its end-to-end system design. QOS’s components synergistically optimize fidelity and resource efficiency while supporting composable techniques within individual layers.
  3. 3.
    QOS systematically explores key tradeoffs in quantum computing: Fidelity versus overheads in error mitigation and performance estimation, as well as fidelity versus QPU utilization and waiting times in multi-programming and scheduling, respectively.
  4. 4.
    QOS introduces new ideas across the quantum software stack. These include the Qernel abstraction as a unifying execution unit, the composable error mitigation pipeline combining complementary techniques, a multi-programming model that quantifies the compatibility and effective utilization of co-located jobs, and the first fidelity-aware multi-objective scheduler.
Figure 2: Technical Foundations (§ 2.2) (a) The quantum circuit of Figure 1. (b) The physical layout of an IBM Falcon QPU. (c) The transpiled circuit with the QPU’s noise sources.
2 Background
2.1 Quantum Computing 101
Quantum computers solve specific computationally hard problems exponentially faster than classical computers by leveraging the quantum mechanics principles of superposition and entanglement. Specifically, the basic units of quantum computation are qubits, which during quantum computation are both 0 and 1 at the same time (recall Schrodinger’s cat experiment [65]), and when entangled, they can interact with each other even over large distances [41].
To solve NP-hard problems (e.g., maximum-cut) with quantum computers, we use algorithms such as the Quantum Approximate Optimization Algorithm (QAOA) [20]. This algorithm is considered practical for today’s quantum hardware capabilities and influential towards achieving quantum advantage [63].
Foundational example. Figure 1 shows how QAOA solves an example max-cut problem for an input graph (a). First, the graph is encoded as a quantum circuit. Quantum circuits comprise qubits and quantum gates, akin to logical gates in classical circuits (e.g. NOT, XOR), which are applied over time (from left to right). In the max-cut case, each graph vertex corresponds to a qubit and each edge to a 2-qubit gate in the circuit (b). At the end of the circuit, we measure each qubit to read its value, which gives bitstrings as output. Notably, measurements collapse the superposition state to a definite binary value.
@@ 2.1 Quantum Computing 101
Figure 3: (a) Challenge #1, Fidelity (§ 3.1). Impact of the number of qubits (circuit size) on fidelity. There is an average 98.9% reduction in fidelity from 4 to 24 qubits. (b) Challenge #2, Spatial heterogeneity (§ 3.2). Fidelity of a 12-qubit GHZ circuit on different IBM QPUs. There is a 38% fidelity difference from best to worst QPU.
Since quantum mechanics is inherently probabilistic, the bitstring we get is truly random, but due to the algorithm structure, it follows a probability distribution, which ultimately renders quantum computers useful. Thus, we execute the circuit in many trials (“shots”), with each trial providing a specific bitstring from the qubit measurements. The solution of the quantum calculation is, therefore, a probability distribution over all possible bitstrings of the measured qubits, where high probability maps to the solution, while low ($\sim 0$) does not represent a solution (c). In our example, the bitstring 10001 represents the max-cut solution that contains the partitions {a, d} and {b, c, e} (d).
2.2 Technical Foundations
@@
Execution Model. The technology and engineering required to build QPUs renders them an expensive resource, i.e., there are less than 100 QPUs globally offered in the cloud in a quantum-as-a-service fashion [32, 5, 24, 3]. To run quantum programs, users typically write circuit-level code (Figure 2 (a)), which then transpile on the QPU to make it executable, send it to the cloud for execution, and finally get the results back. Specifically, the transpilation process performs three key steps: (1) converting the gates of the circuit to the native gate set of the QPU, (2) mapping the logical qubits of the circuit to the physical qubits of the QPU, (3) routing the qubits to the physical qubits with restrictive connectivity by inserting additional costly gates. Figure 2 (b) shows the physical layout of an IBM Falcon QPU. Vertices are the physical qubits, and the edges capture their connectivity, i.e., between which qubits we can apply 2-qubit gates. Figure 2 (c) shows the physical circuit after transpilation with the QPU’s noise characteristics, which we detail next.
@@
QPU characteristics. Today’s QPUs are described as noisy intermediate-scale quantum (NISQ) devices [57] since they exhibit low qubit numbers (e.g., up to a few 100s [32]) and are susceptible to hardware and environmental noise. Specifically, when measuring a qubit, there is a chance to read the opposite value, and when applying gates, there is a chance the gate performs a wrong operation [23]. On top of that, when qubits are left idle (no gates applied) for more than a few hundred microseconds, the superposition decoheres to the |$\ket{0}$ state [35], similar to resetting a register to 00. Lastly, qubits destructively interfere with each other via crosstalk effects [45]. Figure 2 (c) shows qubits $Q_{0}$ and $Q_{3}$ that influence each other via crosstalk, noisy gates, qubit $Q_{0}$ that is left idle for long enough to decohere, and noisy measurements.
@@
QPU heterogeneity. Additionally, QPUs are vastly heterogeneous across space and time, unlike classical accelerators. Across space, QPUs vary in terms of technology, e.g., superconducting qubits [69, 19] or trapped ions [30], architectures of the same technology, e.g., Falcon or Osprey superconducting QPUs [32], and noise properties (formally called noise model) even for the same architecture [23], e.g., two identical QPUs exhibit different noise errors, etc. Across time, the QPUs are calibrated regularly to maintain their performance [82, 77, 31], a process that generates calibration data. These data quantify the noise errors and change unpredictably after each calibration cycle.
@@
Performance metric. To measure the quality of a circuit execution on NISQ QPUs, we use the fidelity metric [21], which measures the similarity between the noisy probability distribution and the ideal probability distribution that noiseless, ideal QPUs can obtain. Given two probability distributions over measurement outcomes, the ideal distribution $P_{ideal}$ and the noisy distribution $P_{noisy}$, fidelity is computed as: $F(P_{ideal},P_{noisy})=(\sum_{i}\sqrt{P_{ideal}(i)*P_{noisy}(i)})^{2}$. Fidelity is a number in the $[0,1]$ range, where a higher fidelity means a better quality result.
Circuit properties. Execution fidelity is impacted by the aforementioned QPU noise model and the circuit properties that describe the circuit’s complexity and, thus, susceptibility to errors. Circuit width refers to the number of logical qubits involved in the circuit and typically also the physical (H/W) qubits required to execute it. Circuit depth denotes the largest number of layers of gates applied to the qubits and quantifies the circuit runtime. A larger depth typically indicates a higher chance for decoherence errors. Last, the number of non-local gates refers to the gates that act on non-adjacent qubits, which are particularly susceptible to errors in NISQ devices. Generally, larger circuits (larger width, depth, or number of non-local gates) equals lower fidelity.
Figure 4: (a) Challenge #2, Temporal variance (§ 3.2). Fidelity of a 6-qubit GHZ circuit on IBM Perth, across 120 calibration days. There are 20 pairs of days with more than 5% difference in fidelity. (b) Challenge #3, Utilization (§ 3.3). Maximum utilization achieved on a 27-qubit QPU for nine benchmarks while maintaining at least 0.75 fidelity. The average utilization is $26.3\%$, and the max is $29.6\%$. (c) Challenge #4, QPU Load (§ 3.4). Number of pending jobs on different IBM QPUs. The groups separated by vertical red lines indicate QPUs of the same size. There is up to 57$\times$ difference in number of jobs between QPUs of the same size.
Circuit feature vectors. While circuit properties are useful, sometimes they fail to characterize the structural and computational properties of quantum circuits. Thus, we can use the Supermarq feature vectors, a set of six metrics proposed by Tomesh et al. [76] with the initial goal of quantifying quantum benchmark coverage. These features quantify the parallelism and speedup potential, the lower bound on execution time, QPU connectivity requirements, entanglement (i.e. complexity), and susceptibility to decoherence and measurement errors.
3 Motivation and Key Ideas